# url]sites.google.com/view/mslee2

### v1

@@ 3 Physics-Informed Neural Networks for the P2D model
$+$ $\left|D_{s}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{s}}{\partial x}(t,L_{p}+L_{s})-D_{n}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{n}}{\partial x}(t,L_{p}+L_{s})\right|^{2}+\left|\widetilde{c}_{s}(t,L_{p}+L_{s})-\widetilde{c}_{n}(t,L_{p}+L_{s})\right|^{2}dt$
The boundary, initial, and interface losses for other variables are defined similarly, following the same manners. This conventional approach is illustrated in Figure 1.
Figure 1: A schematic diagram of the PINN architecture for the P2D model of a lithium-ion battery cell.
While PINNs, in principle, can be applied to solve a wide range of PDEs, they have been known to exhibit failure modes [32], particularly in complex non-linear PDEs. When applying the vanilla PINN approach to the P2D model, it often encounters such failure modes due to the system’s inherent complexity and nonlinearity. In the remainder of this section, we introduce methods tailored specifically for the P2D model to improve PINNs’ performance.
3.1 Multi-Scale Issues
One of the challenges in applying PINNs to the P2D model lies in the significant scale differences between the microscopic and macroscopic coordinates, as well as the various physical and chemical constants in the system. The P2D model considers two spatial coordinates: (1) macroscopic coordinates $x$, representing the through-thickness position across the electrolyte and electrodes, and (2) microscopic coordinates $r$, representing the radial position within solid particles. These scale differences pose numerical difficulties in the optimization process, particularly when the scales differ by several orders of magnitude. Additionally, the physical and chemical constants, such as diffusion coefficients, and conductivity, exhibit vast disparities in scale, further complicating the learning process in the PINN framework.
Such scale-related issues have been widely reported not only in classical numerical methods but also in several SciML methods. One of the simplest but effective techniques to address these issues is non-dimensionalization. Following this, we adopt a non-dimensionalized form of the P2D model to mitigate the scaling problems. Specifically, the coordinate components $t$, $x$ and $r$, and the state variable $c_{s,i}$, $c_{i}$, $\phi_{i}$, and $\psi_{i}$ are rescaled as follows:
$\hat{t}=\frac{t}{\mathcal{T}},\ \hat{x}=\frac{x}{L},\ \hat{r}=\frac{r}{\mathcal{R}_{i}},\ \hat{c}_{s,i}=\frac{c_{s,i}}{c_{s,\mathrm{max}}},\ \hat{c}_{i}=\frac{c_{i}}{c_{\mathrm{ref}}},\ \hat{\phi_{i}}=\frac{\phi_{i}}{\phi_{\mathrm{ref}}},\ \text{and}\ \hat{\psi}_{i}=\frac{\psi_{i}}{\psi_{\mathrm{ref}}},$
which gives $\hat{t},\ \hat{x},\ \hat{r}\in[0,1]$, $\hat{c}_{s,i}\in[0,1]$, and $\hat{c}_{i}$, $\hat{\phi}_{i}$, $\hat{\psi}_{i}\sim\mathcal{O}(1)$. Following this non-dimensionalization, the rescaled governing equations are given as follows:
@@ 3.1 Multi-Scale Issues
Butler-Volmer Equation With the rescaled state variables $\hat{c}_{s,i}$, $\hat{c}_{i}$, $\hat{\phi}_{i}$, and $\hat{\psi}_{i}$, the formula for the reaction rate $j_{i}$ is also rewritten as follows:
$j_{i}=\begin{cases}2k_{i}c_{s,i,\text{max}}c_{\text{ref}}^{0.5}\left(1-\hat{c}_{s,i,\mathrm{surf}}\right)^{0.5}\hat{c}_{s,i,\mathrm{surf}}^{0.5}\hat{c}_{i}^{0.5}\sinh\left(\dfrac{F}{2RT}\eta_{i}\right)&\text{for}\ i=p,n\\ 0&\text{for}\ i=s\end{cases}$
@@
3.2 Initial and Boundary conditions
@@
To address this issue, an alternative approach is to redesign the neural network architecture so that it inherently satisfies the boundary and initial conditions, eliminating the need for additional penalty loss terms. This method, commonly known as the hard constraint method, has been widely applied in PINN applications [34]. Adopting this hard constraint approach, we reconstruct the network architectures in such a way that it automatically satisfies some of the boundary and initial conditions, as follows:
@@
Through this hard-constraint method, the approximate solutions by the neural networks automatically satisfy the initial conditions and Dirichlet boundary conditions without the need for imposing the initial and Dirichlet boundary conditions through soft constraints. However, for Neumann boundary conditions, we do not apply the hard-constraint method and instead continue using the penalty method for training.
3.3 Interfacial conditions
@@
To explore how PINNs handle complex interface conditions, we tested a simplified paradigm example to represent the geometry of a lithium-ion battery:
@@
This paradigm example represents a simplified scenario that reflects the essential dynamics of a lithium-ion battery system. It enables us to validate the performance of the domain decomposition method in handling interfacial conditions with discontinuous derivatives. By applying appropriate boundary and interface conditions, this setup captures the behavior of the electrodes and the separator.
To solve this system using PINNs, we use five neural networks—one for each variable. Interface conditions are incorporated into the physics-informed loss function by adding appropriate penalties.
@@
Figure 2 shows the result of the simulation. Despite the complexity of the system, which includes four boundary conditions, six interface conditions, and five PDEs (5), the PINN framework successfully approximates the exact solution with only the penalty method on the interfacial conditions. Through this paradigm example, we found that a simpler approach—adding soft penalties—is surprisingly sufficient to ensure proper information exchange between the subdomains.
In this section, we have addressed multi-scale issues and the enforcement of boundary, initial, and interfacial conditions in applying PINNs to the P2D model. While these general techniques help mitigate many common challenges, more sophisticated approaches are still required to effectively solve the full P2D model. The most significant challenge arises from the highly non-linear Butler-Volmer equation. In the following section, we will discuss the complexities posed by the BV equations of the P2D mdoel in detail.
4 Highly non-linear Butler-Volmer equation
The Butler-Volmer (BV) equation (1), which governs the electrochemical reaction rates at the interface between the solid particles and the electrolyte in the P2D model, poses major challenges in applying Physics-Informed Neural Networks (PINNs). The BV equation is highly nonlinear, making it one of the most challenging aspects of the P2D model to address through PINN frameworks.
To better understand this, consider the rescaled governing equation for the solid potential $\psi_{p}$:
$\frac{\partial^{2}\psi_{p}}{\partial x^{2}}=\frac{L^{2}a_{p}Fj_{p}}{\sigma_{p}^{\mathrm{eff}}\psi_{\mathrm{ref}}},\qquad\text{for}\ x\in(0,1),$
where the source term on the right-hand side is given as:
@@ 4 Highly non-linear Butler-Volmer equation
$\underbrace{\frac{2kc_{s,p,\text{max}}c_{\text{ref}}^{0.5}L^{2}a_{p}F}{\sigma^{\mathrm{eff}}\psi_{\mathrm{ref}}}\left(1-c_{s,p,\mathrm{surf}}\right)^{0.5}c_{s,p,\mathrm{surf}}^{0.5}c_{p}^{0.5}}_{=:\mathcal{H}}\sinh\bigg{(}\underbrace{\frac{F}{2RT}}_{=:\mathcal{G}}\eta_{p}\bigg{)},$ (7)
@@
according to the definition of $j_{p}$ in (1). Notably, the constant $\mathcal{G}$ inside the hyperbolic sine function in (7) is approximately 19.5. As shown in the left plot of Figure 3, the graph of $f(\eta)=\sinh(19.5\eta)$ demonstrates a very fast growth rate as $\eta$ increases. Recalling that $\eta_{p}=\phi_{p}-\psi_{p}-U_{p}$, we note that small changes in the approximate solutions $\phi_{p}$ and $\psi_{p}$ can lead to large fluctuations in the loss function, exacerbating instability during training. The impact of this issue is illustrated in the upper right panel of Figure 3, which shows the loss behavior obtained using the conventional PINN approach described earlier. During the early stages of training, the initial loss reaches a magnitude of approximately $O(10^{75})$. In contrast, when $\eta$ is manually set to zero, the starting loss value drops to approximately $O(10^{-1})$, clearly demonstrating the ill-conditioning posed by the BV equation.
@@
Figure 3: The left panel displays the graph of the function $f(\eta)=\sinh(19.5\eta)$, while the right panel illustrates how errors in $\eta$ are amplified when evaluating the PINN loss. The solid line represents the baseline, whereas the dashed line corresponds to the scenario where $\eta=0$ is manually enforced. These results highlight why the nonlinear Butler-Volmer equation (1) is the primary contributor to the large PINN loss values.
On the other hand, a different issue also arises. The term $\mathcal{H}$ in (7) is relatively smaller in magnitude compared to other variables, which are rescaled to be on the order of $1$. Although the $\sinh$ function grows exponentially, it behaves linearly near the origin, where the over-potential $\eta$ lies in most practical scenarios. Due to the smallness of $\mathcal{H}$, the source term (7) exhibits only very small linear changes when $\eta\ll 1$. This implies that even if the approximate solution $\widetilde{\psi}_{p}$ does not match the true solution $\psi_{p}$ and is instead shifted by a constant, the shifted function $\widetilde{\psi}_{p}$ can still yield a relatively small PDE residual loss. Furthermore, the boundary loss would remain exactly zero. As a result, it is necessary to capture these small variations at a highly precise level to distinguish between $\psi_{p}$ and $\widetilde{\psi}_{p}$. This issue complicates the application of PINNs to solving the P2D model.
In order to illustrate the practical impact of this issue posed by the nonlinear BV equation, we consider a simplified paradigm example. This example is designed to experimentally showcase how the issue by a small change of the source term could make a significant challenge when applying PINNs to the P2D model, emphasizing the importance of addressing these challenges effectively. The paradigm example is given as follows:
@@
$U(x)$ $=x^{2}(x-1.5)+4-10^{-5}(1-x)^{10}+\log\left(j_{0}^{-1}(10^{-3}-10^{-4})(1-x)^{8}+1\right),$
and $j_{0}$ is a parameter to be chosen later. The exact solutions for $\phi$ and $\psi$ are:
$\phi(x)$ $=10^{-5}(1-x)^{10},\qquad\text{and}\qquad\psi(x)=-x^{2}(x-1.5)-4.$
We note that in this problem, the term $j_{0}$ plays a role similar to $\mathcal{H}$ in the original P2D model.
Let us now consider a shifted function $\widetilde{\psi}$ given as $\widetilde{\psi}=\psi+\epsilon$, then by using Taylor expansion:
$e^{\widetilde{\psi}-\phi+U}-1\approx e^{\psi-\phi+U}-1+e^{\psi-\phi+U}\epsilon=\dfrac{j}{j_{0}}+e^{\psi-\phi+U}\epsilon.$
Letting $\widetilde{j}=j_{0}\left(e^{\widetilde{\psi}-\phi+U}-1\right)$ , we obtain:
$\widetilde{j}\approx j+j_{0}e^{\psi-\phi+U}\epsilon.$
Since $e^{\psi-\phi+U}=j_{0}^{-1}(10^{-3}-10^{-4})(1-x)^{8}+1\leq j_{0}^{-1}(10^{-3}-10^{-4})+1$, the shifted function $\widetilde{\psi}$ results in a change in the source term $j$ by approximately $j_{0}\epsilon+10^{-3}\epsilon$. Therefore, the PDE residual of the shifted function $\widetilde{\psi}$ is
$|\widetilde{\psi}_{xx}+f-\widetilde{j}|$ $=|j-\widetilde{j}|\qquad(\text{since}\ \widetilde{\psi}_{xx}=\psi_{xx})$
$=O(j_{0}\epsilon+10^{-3}\epsilon),$
while the boundary loss is exactly zero as the boundary conditions of $\psi$ are Neuman-type. When $\epsilon=O(1)$ and $j_{0}=10^{-1}$, a PDE residual loss of $O(10^{-2})$ from $\widetilde{\psi}$ suggests that the shift has a significant impact. However, if $j_{0}$ is much smaller, such as $j_{0}=10^{-3}$, the PDE residual loss would need to be below $O(10^{-6})$ to capture the constant shift.
In typical PINN training, it is generally impossible to minimize the loss function exactly to zero, and the loss typically converges to an order of $10^{-4}$ or $10^{-5}$. However, where the required PDE residual loss for $\widetilde{\psi}$ must be smaller than the achieved loss, the PINN may be unable to distinguish between the true solution $\psi$ and the slightly shifted function $\widetilde{\psi}$ as both result in nearly identical loss values at that precision. Consequently, when $j_{0}$ is very small, the network struggles to accurately identify the true solution.
The experimental results presented in Figure 4 support our observation. Specifically, for a relatively large $j_{0}=10^{-1}$, the PINN can find the correct solution with ease. However, for a smaller $j_{0}=10^{-3}$, the PINN solution is incorrect, which looks shifted from the true solution. This aligns with our observation that PINNs struggle to distinguish between the true solution and a solution shifted by a constant when $j_{0}$ is small. PINNs fail to differentiate small changes in the PDE residuals, leading to an inaccurate prediction.
Figure 4: PINN results for the paradigm model (8) with $j_{0}\in\{10^{-1},10^{-2},10^{-3}\}$. The vertical axes represent absolute error.
This paradigm example illustrates a key challenge that comes from the highly nonlinear nature of the BV equation, combined with small scaling factors $\mathcal{H}$, which makes it difficult for the networks to accurately capture the true solution. In conclusion, this observation emphasizes the importance of carefully addressing the issue from the BV equation when applying PINNs to the P2D model.
@@
It is worth highlighting the study by Hassanaly et al. [35], which faced similar challenges when training PINNs for the SPM with the nonlinear BV equation. Their solution involved a two-step training approach: first, training on a linearized version of the BV equation and then transferring the learned parameters to the nonlinear model as initialization. While innovative, this method comes with certain limitations–it introduces additional hyperparameters that require careful tuning, and detailed results of their approach were not provided. In contrast, we focus on developing an end-to-end training strategy that avoids these complexities, offering a simpler and more robust alternative for tackling the nonlinear BV equation of the P2D model.
5 Strategies for the non-linear Butler-Volmer equation
As discussed in the previous sections, the highly nonlinear nature of the Butler-Volmer (BV) equation can cause significant difficulties in applying PINNs to solve the P2D model. In this section, we propose two strategies aimed at mitigating these challenges: the introduction of a bypassing term to stabilize the training process and the addition of secondary conservation laws to improve the accuracy of solutions.

### v2

@@ 3.1 Vanilla PINNs for the P2D model
$+$ $\left|D_{s}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{s}}{\partial x}(t,L_{p}+L_{s})-D_{n}^{\mathrm{eff}}\frac{\partial\widetilde{c}_{n}}{\partial x}(t,L_{p}+L_{s})\right|^{2}+\left|\widetilde{c}_{s}(t,L_{p}+L_{s})-\widetilde{c}_{n}(t,L_{p}+L_{s})\right|^{2}\mathrm{d}t.$
The boundary, initial, and interface losses for other variables are defined similarly, following the same manners. While the vanilla PINN approach, in principle, be applied to a wide range of PDEs, PINNs have often been shown to exhibit failure modes [32], particularly for strongly nonlinear PDE systems. Indeed, when applying this straightforward method to the P2D model, one encounters such failure modes due to inherent complexities. In the remainder of this section, we discuss several challenges with the vanilla approach and address them by using standard strategies.
3.2 Issues with vanilla PINNs
This subsection is devoted to addressing several issues that arise when applying the above vanilla approach to solve the P2D model.
Multi-scale issues
One of the challenges in applying PINNs to the P2D model lies in the significant scale differences between the microscopic and macroscopic coordinates, as well as the various physical and chemical constants in the system. The P2D model considers two spatial coordinates: (1) macroscopic coordinates $x$, representing the through-thickness position across the electrolyte and electrodes, and (2) microscopic coordinates $r$, representing the radial position within solid particles. These scale differences pose numerical difficulties in the optimization process, particularly when the scales differ by several orders of magnitude. Additionally, the physical and chemical constants, such as diffusion coefficients and conductivity, exhibit vast disparities in scale, further complicating the learning process in the PINN framework.
Such scale-related issues have been widely reported not only in classical numerical methods but also in several SciML methods. One of the simplest yet effective techniques to address these issues is non-dimensionalization. Following this, we adopt a non-dimensionalized form of the P2D model to mitigate the multi-scale issues. Specifically, the coordinate components $t$, $x$, and $r$, and the state variables $c_{s,i}$, $c_{i}$, $\phi_{i}$, and $\psi_{i}$ are rescaled as follows:
$\hat{t}=\frac{t}{\mathcal{T}},\quad\hat{x}=\frac{x}{L},\quad\hat{r}=\frac{r}{\mathcal{R}_{i}},\quad\hat{c}_{s,i}=\frac{c_{s,i}}{c_{s,\mathrm{max}}},\quad\hat{c}_{i}=\frac{c_{i}}{c_{\mathrm{ref}}},\quad\hat{\phi_{i}}=\frac{\phi_{i}}{\phi_{\mathrm{ref}}},\quad\hat{\psi}_{i}=\frac{\psi_{i}}{\psi_{\mathrm{ref}}},$
which yields $\hat{t},\ \hat{x},\ \hat{r}\in[0,1]$, $\hat{c}_{s,i}\in[0,1]$, and $\hat{c}_{i}$, $\hat{\phi}_{i}$, $\hat{\psi}_{i}\sim\mathcal{O}(1)$. Following this non-dimensionalization, the rescaled governing equations are given by:
@@ 3.2 Issues with vanilla PINNs
Butler–Volmer equation With the rescaled state variables $\hat{c}_{s,i}$, $\hat{c}_{i}$, $\hat{\phi}_{i}$, and $\hat{\psi}_{i}$, the formula for the reaction rate $j_{i}$ is also rewritten as follows:
$j_{i}=\begin{cases}2kc_{s,i,\text{max}}c_{\text{ref}}^{0.5}\left(1-\hat{c}_{s,i,\mathrm{surf}}\right)^{0.5}\hat{c}_{s,i,\mathrm{surf}}^{0.5}\hat{c}_{i}^{0.5}\sinh\left(\dfrac{F}{2RT}\eta_{i}\right)&\text{for}\ i=p,n\\ 0&\text{for}\ i=s\end{cases}$
@@
Initial and boundary conditions
@@
To address this issue, an alternative approach is to redesign the neural network architecture so that it inherently satisfies the boundary and initial conditions, eliminating the need for additional penalty loss terms. This method, commonly known as the hard constraint method, has been widely applied in PINN applications [34]. Adopting this hard-constraint approach, we reconstruct the network architectures in such a way that it automatically satisfies some of the boundary and initial conditions, as follows:
@@
Through this hard-constraint method, the approximate solutions automatically satisfy the initial conditions and Dirichlet boundary conditions without soft constraints. However, for Neumann boundary conditions, we do not apply the hard-constraint method and instead continue using penalty methods, to be consistent with interfacial conditions which will be discussed below.
Interfacial conditions
@@
To explore how PINNs handle complex interface conditions, we tested a simplified toy example to represent the geometry of a lithium-ion battery:
@@
This toy example represents a simplified scenario that reflects the essential dynamics of a lithium-ion battery system. It enables us to validate the performance of the domain decomposition method in handling interfacial conditions with discontinuous derivatives. By applying appropriate boundary and interface conditions, this setup captures the behavior of the electrodes and the separator.
To solve this system using PINNs, we use five neural networks—one for each variable. Interface conditions are incorporated into the PINN loss by adding appropriate penalties.
@@
Figure 2 shows the result of the simulation. Despite the complexity of the system, which includes four boundary conditions, six interface conditions, and five PDEs (5), the PINN framework successfully approximates the exact solution with only the penalty method on the interfacial conditions. Through this toy example, we found that a simpler approach—adding soft penalties—is surprisingly sufficient to ensure proper information exchange between the subdomains.
3.3 Limitations of the conventional approach
In this section, we presented a vanilla PINN approach for solving the P2D model and then applied existing strategies to address certain issues with that approach. The resulting conventional PINN framework is illustrated in Figure 3. Notably, all initial conditions are already satisfied by construction through the hard constraint method, making any additional penalty terms for initial conditions unnecessary. However, while these strategies mitigate several difficulties that may arise in solving the P2D model, they remain insufficient for fully resolving complexities of the P2D model. In particular, when using a conventional PINN setup, the initial training loss can be as large as $O(10^{75})$, leading to severe instability and ultimately causing convergence to fail. In the following section, we focus on the challenges posed by the BV equation, which we believe is a key reason why this conventional approach can fail.
Figure 3: A schematic diagram of the PINN architecture for the P2D model of a Li-ion battery cell.
4 Challenges related to the Butler–Volmer equation
In this section, we examine the remaining issues of the conventional PINN approach in Section 3, which lead to an excessively large final loss and ultimately fail to yield accurate solutions. We have identified that these problems stem primarily from the nonlinearity of the BV equation (1) in the P2D model. This strong nonlinearity makes the BV equation one of the most challenging parts of the P2D model to address using PINNs.
4.1 Sensitivity of the BV equation with respect to $\eta$
We recall the BV equation in its rescaled form:
@@ 4.1 Sensitivity of the BV equation with respect to $\eta$
$j_{i}\;=\;2k\,c_{s,i,\text{max}}\,c_{\text{ref}}^{0.5}\,\bigl{(}1-c_{s,i,\mathrm{surf}}\bigr{)}^{0.5}\,c_{s,i,\mathrm{surf}}^{0.5}\,c_{i}^{0.5}\,\sinh\!\bigg{(}\underbrace{\frac{F}{2RT}}_{\mathcal{G}}\,\eta_{i}\bigg{)},$ (7)
@@
where $\eta_{i}=\psi_{i}-\phi_{i}-U_{i}$. Notably, the constant $\mathcal{G}$ inside the hyperbolic sine function in (7) is approximately 19.5. As shown in the left plot of Figure 4, the graph of $f(\eta)=\sinh(19.5\eta)$ demonstrates a very fast growth rate as $\eta$ increases. We note that small changes in the approximate solutions $\phi_{p}$ and $\psi_{p}$ can lead to large fluctuations in the loss function, exacerbating instability during training. The impact of this issue is illustrated in the upper right panel of Figure 4, which shows the loss behavior obtained using the conventional PINN approach described earlier. During the early stages of training, the initial loss reaches a magnitude of approximately $O(10^{75})$. In contrast, when $\eta$ is manually set to zero, the starting loss value drops to approximately $O(10^{-1})$, clearly demonstrating the ill-conditioning posed by the BV equation.
@@
Figure 4: The left panel displays the graph of the function $f(\eta)=\sinh(19.5\eta)$, while the right panel illustrates how errors in $\eta$ are amplified when evaluating the PINN loss. The solid line represents the baseline, whereas the dashed line corresponds to the scenario where $\eta=0$ is manually enforced. These results highlight why the nonlinear BV equation (1) is the primary contributor to the large PINN loss values.
4.2 Discrepancy between loss and accuracy
A small PINN loss does not always imply that the corresponding approximate solution is sufficiently close to the true solution of the target PDEs. Such discrepancies between loss and accuracy have been reported in the literature [32, 35]. In the context of the P2D model, this issue is especially relevant for the BV equation. We recall the rescaled governing equation for the solid potential $\psi_{p}$:
$\frac{\partial^{2}\psi_{p}}{\partial x^{2}}\;=\;\frac{L^{2}a_{p}F\,j_{p}}{\sigma_{p}^{\mathrm{eff}}\,\psi_{\mathrm{ref}}},\quad x\in(0,1),$
where the source term on the right-hand side is:
$\underbrace{\frac{2k\,c_{s,p,\max}\,c_{\text{ref}}^{0.5}\,L^{2}\,a_{p}\,F}{\sigma^{\mathrm{eff}}\,\psi_{\mathrm{ref}}}\,\bigl{(}1-c_{s,p,\mathrm{surf}}\bigr{)}^{0.5}\,c_{s,p,\mathrm{surf}}^{0.5}\,c_{p}^{0.5}}_{=:\mathcal{H}}\;\sinh\Bigl{(}\tfrac{F}{2RT}\,\eta_{p}\Bigr{)},$ (8)
according to the definition of $j_{p}$ in (1). Notably, the coefficient $\mathcal{H}$ can be much smaller in magnitude than other rescaled variables (on the order of 1). Although the hyperbolic sine function grows exponentially for large inputs, it behaves almost linearly near the origin. In most practical situations, the over-potential $\eta$ remains small, so $\sinh\bigl{(}\frac{F}{2RT}\,\eta_{p}\bigr{)}\approx\frac{F}{2RT}\,\eta_{p}$. Consequently, when $\mathcal{H}$ is also very small, the right-hand side of (8) changes only slightly even if $\eta_{p}$ (or $\psi_{p}$) is perturbed.
Keeping these points in mind, let us consider a simple example of how inaccurate solutions can still yield small PINN losses. Suppose $\widetilde{\psi}_{p}$ differs from the true solution $\psi_{p}$ by a constant shift $C=O(1)$, i.e., $\widetilde{\psi}_{p}(x)=\psi_{p}(x)+C$. Then, $\widetilde{\psi}_{p}$ can produce nearly the same small loss as $\psi_{p}$, even though $\widetilde{\psi}_{p}$ is far from the true solution. The reasons are:
  • 1.
    PDE residual: The left-hand side of PDE involves only $\tfrac{\partial^{2}\psi_{p}}{\partial x^{2}}$ so that a constant shift $C$ disappears in that side. Also, if $\mathcal{H}$ is sufficiently small, the source term changes very slightly when shifting $\psi_{p}$ by $C$. Therefore, the shifted one may still lead to a very small PDE residual loss.
  • 2.
    Neumann boundary conditions: The relevant boundary conditions are Neumann conditions so that a constant shift does not change the boundary loss.
An inaccuracy in $\psi_{p}$ can distort other state variables and degrade the solution’s overall accuracy, yet the PINN loss function may fail to detect it. In other words, even if the PINN loss is minimized to a very small value, the approximate solutions may still deviate substantially from the true solution. This discrepancy between loss and accuracy complicates the application of PINNs to the P2D model, underscoring the need for additional strategies to mitigate it.
In addition, we present the following toy example to illustrate the practical impact of how the small coefficient $\mathcal{H}$ affects the accuracy of PINN approaches in a simpler setting:
@@ 4.2 Discrepancy between loss and accuracy
$U(x)$ $=x^{2}(x-1.5)+4-10^{-5}(1-x)^{10}+\log\Bigl{(}j_{0}^{-1}(10^{-3}-10^{-4})(1-x)^{8}+1\Bigr{)},$
and $j_{0}$ is a parameter to be chosen. The exact solutions for $\phi$ and $\psi$ are:
$\phi(x)=10^{-5}(1-x)^{10},\quad\psi(x)=-x^{2}(x-1.5)-4.$
Here, $j_{0}$ plays a role similar to $\mathcal{H}$ in the original P2D model.
Let us now consider a shifted function $\widetilde{\psi}=\psi+C$. Using a Taylor expansion, we have:
$e^{\widetilde{\psi}-\phi+U}-1\;\approx\;e^{\psi-\phi+U}-1+C\,e^{\psi-\phi+U}\;=\;\frac{j}{j_{0}}+C\,e^{\psi-\phi+U}.$
Let $\widetilde{j}=j_{0}\bigl{(}e^{\widetilde{\psi}-\phi+U}-1\bigr{)}$. Then
$\widetilde{j}\approx j+C\,j_{0}\,e^{\psi-\phi+U}.$
Since
$e^{\psi-\phi+U}\;=\;j_{0}^{-1}(10^{-3}-10^{-4})(1-x)^{8}+1\;\leq\;j_{0}^{-1}(10^{-3}-10^{-4})+1,$
the shifted function $\widetilde{\psi}$ alters the source term $j$ by approximately $j_{0}C+10^{-3}C$. Hence, the PDE residual of $\widetilde{\psi}$ becomes
$|\widetilde{\psi}_{xx}+f-\widetilde{j}|$ $=|j-\widetilde{j}|\quad(\text{since }\widetilde{\psi}_{xx}=\psi_{xx}),$
$=O\bigl{(}j_{0}C+10^{-3}C\bigr{)},$
while the boundary loss remains zero (the boundary conditions for $\psi$ are Neumann). According to this observation, when $j_{0}=10^{-1}$, the PDE residual loss from $\widetilde{\psi}$ is on the order of $O(10^{-2})$. However, if $j_{0}$ is much smaller (e.g., $10^{-3}$), the PDE residual would be below $O(10^{-6})$. In typical PINN training, it is generally infeasible to reduce the loss function exactly to zero; the loss typically converges to about $10^{-4}$ or $10^{-5}$. Therefore, if the PINN loss by $\widetilde{\psi}$ is smaller than the achieved loss threshold, the PINN may be unable to distinguish between the true solution $\psi$ and the shifted function $\widetilde{\psi}$. In fact, the experimental results in Figure 5 confirm our observation. Specifically, for a relatively large $j_{0}=10^{-1}$, the PINN provides the correct solution easily, but for a smaller $j_{0}=10^{-3}$, the trained solution is inaccurate. This confirms that the smaller $j_{0}$ is, the more PINNs struggle to distinguish between the true solution and a shifted one.
Figure 5: PINN results for the toy model (9) with $j_{0}\in\{10^{-1},10^{-2},10^{-3}\}$. The vertical axes represent absolute error.
This toy example illustrates a key challenge related to the smallness of $\mathcal{H}$ of the BV equation, which hinder the PINN’s ability to accurately capture the true solution.
@@
It is worth highlighting the study by Hassanaly et al. [36], which encountered similar issues when training PINNs for the single-particle model (SPM) under a nonlinear BV equation. Their solution involved a two-step training approach—first training on a linearized BV equation, then transferring the learned parameters to the nonlinear model for initialization. While innovative, that method introduces extra hyperparameters requiring careful tuning, and their results are not extensively detailed. In contrast, we focus on developing an end-to-end training strategy that avoids these additional complexities, offering a simpler and more robust alternative for handling the nonlinear BV equation in the P2D model.
5 Strategies for the nonlinear Butler–Volmer equation
As discussed in the previous sections, the highly nonlinear nature of the BV equation can cause significant difficulties in applying PINNs to solve the P2D model. In this section, we propose two strategies aimed at mitigating these challenges: the introduction of a bypassing term to stabilize the training process and the addition of secondary conservation laws to improve the accuracy of solutions.